OpenStudy (vera_ewing):

What cosine function represents an amplitude of 3, a period of π, no horizontal shift, and a vertical shift of 2?

OpenStudy (michele_laino):

hint: the amplitude of your function is 2, since we have: \[ - 2 \leqslant 2\cos \left( {\pi x} \right) \leqslant 2\]

OpenStudy (michele_laino):

the requested function is like below: \[y = A\cos \left( {\frac{\pi }{k}} \right) + h\] where A is the amplitude, h is the vertical shift, and k is such taht the period of that function is: \[T = \frac{{2\pi }}{k}\]

OpenStudy (michele_laino):

that is the general formula, for your exercise

OpenStudy (vera_ewing):

So it's going to have to start out as f(x)= 3cos right?

OpenStudy (michele_laino):

yes!

OpenStudy (vera_ewing):

And the last part of the equation will be +2?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

oops.. I have made a typo, here is the right formula: \[y = A\cos \left( {\frac{x}{k}} \right) + h\]

OpenStudy (vera_ewing):

f(x) = 3 cos 2x + 2 ?

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (vera_ewing):

Thank you, Michele! :)

OpenStudy (michele_laino):

thanks! :)